OpenStudy (anonymous):

Please help find the slope of this line...It is fraction so i don' t understand how to do it.

TheSmartOne (thesmartone):

What is the line? :)

OpenStudy (anonymous):

\[G(3, \frac{ 2 }{ 3 }) and H(-\frac{ 2 }{ 3 }, 2)\] I know it is y2-y1 over x2-x1 so \[\frac{ 2 - \frac{ 2 }{ 3 } }{ \frac{ -2 }{ 3 }-3 }\] But, now what?

TheSmartOne (thesmartone):

so correct

TheSmartOne (thesmartone):

\(\sf\Large 2 - \frac{2}{3} = \frac{6}{3} - \frac{2}{3} =?\)

OpenStudy (anonymous):

\[\frac{ 4 }{ 3 }\] Then that would be over \[\frac{ 11 }{ 3 }\] Correct?

TheSmartOne (thesmartone):

correct and now we have \(\sf\Huge \frac{\frac{4}{3}}{\frac{11}{3}} = \frac{4}{3} \div \frac{11}{3} = \frac{4}{3} \times \frac{3}{11}=?\)

OpenStudy (anonymous):

\[\frac{ 12 }{ 33 } ?\]

TheSmartOne (thesmartone):

we can simplify it by dividing by 3 on the numerator and denominator :)

OpenStudy (anonymous):

so 4/11?

TheSmartOne (thesmartone):

correct :)

TheSmartOne (thesmartone):

correct and now we have \(\sf\Huge \frac{4}{3} \times \frac{3}{11}=\frac{4}{\cancel{3}} \times \frac{\cancel{3}}{11}=\boxed{\frac{4}{11}}\)

TheSmartOne (thesmartone):

Good job! :D

OpenStudy (anonymous):

Awesome, Thank you for the help @TheSmartOne it really makes sense now! :)

TheSmartOne (thesmartone):

Anytime! @Savannah_Lynn13 :D